OpenStudy (anonymous):

what is the slope line passing through (-3,1) and (-17,2)

OpenStudy (gudden):

Do you know the formula used to find the slope ?

OpenStudy (anonymous):

m=y2-y1/x2-x1

OpenStudy (anonymous):

m=(2-1)/(-17--3)=1/(-17+3)=1/-14

OpenStudy (gudden):

@angie0520 , well you are new to the site..! Let me make it clear that one is not supposed to give the answer directly in here! Whereas you can help one in going through the question! I insist read the code of conduct

OpenStudy (anonymous):

Oh my bad

OpenStudy (gudden):

BTW, welcome to OS... @angie0520 and @genesis3212 ... And.. be careful next time.... don't feel pity :)

OpenStudy (anonymous):

slope is rise over run. so, find the climb in y and divide by run in x

OpenStudy (gudden):

Co-ordinates are represented as (X, Y) And let me take (-3,1) and (-17,2) points as an example , whose slope we are supposed to find.

OpenStudy (gudden):

 So, -3 is x co-ordinate and y co-ordinate is 1....